Mildred desires to have $7,049 on deposit five years from today. If she has $4,000 to deposit, what rate of interest, compounded annually, must be obtained to accumulate the desired $7,049 in five years? 9% 12% 8% 10%
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

7049=4000*(1+r/100)^5

(7049/4000)^(1/5)=(1+r/100)

(1+r/100)=1.12

r=1.12-1

=12%